OpenStudy (anonymous):

College Algebra: I'm reviewing for my Final coming up and I need tons of help. Some of the math problems I'm going through are not making sense. The center of the circle with the following equation is: (comment below)

OpenStudy (anonymous):

\[(x-4)^{2}+(y+9)^{2}=25\]

OpenStudy (solomonzelman):

you will likely need logarithms, rules for shifts of functions, circles and other like staff....

OpenStudy (solomonzelman):

Have you like made a big outline of things you need to know?

OpenStudy (solomonzelman):

you just need to know about circles?

OpenStudy (solomonzelman):

\(\large\color{black}{ (x-h)^2+(y-k)^2=r^2 }\) where \(\large\color{black}{ (h,k) }\) is the center of the circle.

OpenStudy (anonymous):

Well, I had a standard form of a circle in mind, but then I started thinking it wouldn't fit. It's \[(x _{2}-h)^{2}+(y _{2}-k)^{2}=r ^{2}\]

OpenStudy (anonymous):

Yeah, there you go lolx

OpenStudy (solomonzelman):

it is not x sub of something, just \(\large\color{black}{ (x-h)^2+(y-k)^2=r^2 }\)

OpenStudy (solomonzelman):

Do you have any problems to work with, or want me to think of an example?

OpenStudy (anonymous):

I gave you the equation in the first comment.

OpenStudy (solomonzelman):

Sorry, I was blind.. okay, and what do you want, to extract the center of it?

OpenStudy (anonymous):

I believe so, the math problem only said "The center of the circle with the following equation is:"

OpenStudy (solomonzelman):

\(\large\color{black}{ (x\color{blue}{-h})^2+(y\color{red}{-k})^2=r^2 }\) \(\large\color{black}{ (x\color{blue}{-4})^2+(y\color{red}{+9})^2=r^2 }\) `center is (h,k)` can you compare the equations?

OpenStudy (solomonzelman):

first, can you tell me the h ?

OpenStudy (anonymous):

Basically -4 and 9 is the center of the circle.

OpenStudy (solomonzelman):

no it is the other way,

OpenStudy (anonymous):

So, 4 and -9?

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ (x-\color{blue}{4})^2+(y+9)^2 =25 }\) \(\large\color{black}{ (x-\color{blue}{4})^2+(y-\color{red}{-9})^2 =25 }\) \(\large\color{black}{ (x-\color{blue}{4})^2+(x-\color{red}{h})^2 =r^2 }\)

OpenStudy (solomonzelman):

I am re-writing the y+9, as y-(-9), to give it a form of (y-k). in this case, k is -9. See?

OpenStudy (anonymous):

Yes I do :)

OpenStudy (solomonzelman):

GOod;) If I were to ask you, What is the center of the following circle: \(\large\color{black}{ (x+3)^2+(y+4)^2 =16 }\) what would you tell me?

OpenStudy (anonymous):

That the center of the circle is (-3, -4)

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

Yes, I got it!! :D

OpenStudy (solomonzelman):

And a center of the above? \(\large\color{black}{ x^2+(y-2)^2 =16 }\)

OpenStudy (anonymous):

Oh, that's tricky. isn't it (-2, 256) or (1, -2)?

OpenStudy (solomonzelman):

Hint: \(\large\color{black}{ (x-0)^2+(y-2)^2 =16 }\)

OpenStudy (anonymous):

Oh, (0,2) lolx

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

And lastly, \(\large\color{black}{ x^2+y^2 =0 }\) ?

OpenStudy (solomonzelman):

I am tricking you...

OpenStudy (anonymous):

(0,0) :)

OpenStudy (anonymous):

So i noticed :D

OpenStudy (anonymous):

You're also teaching me, which is good. I do better with repetition :)

OpenStudy (solomonzelman):

No.

OpenStudy (solomonzelman):

When \(\large\color{black}{ x^2+y^2 =0 }\) comparing to \(\large\color{black}{ (x-h)^2+(y-k)^2 =r^2 }\) our radius, the r, is zero. and this is therefore not a circle.

OpenStudy (anonymous):

So, no solution right?

OpenStudy (solomonzelman):

yes, it is not a circle.

OpenStudy (anonymous):

That was a tricky one

OpenStudy (solomonzelman):

okay, and how about this, \(\large\color{black}{ (x-2)^2+(y-2)^2 =-2 }\) ?

OpenStudy (anonymous):

I'm thinking it's (2,2), but that -2 is making me second guess myself.

OpenStudy (solomonzelman):

a radius can;t be an imaginary number, because if r^2=-2, then \(\large\color{black}{ r=2i }\).

OpenStudy (anonymous):

Ohhh. So it's not a circle because of the radius

OpenStudy (solomonzelman):

it is not a circle again. The radius has to be greater than zero.

OpenStudy (solomonzelman):

yes, i's not a circle because of the radius. And lastly, \(\large\color{black}{ (x+1)^2+y^2=9 }\)

OpenStudy (solomonzelman):

I keep saying and lastly and conttinuing... lol

OpenStudy (anonymous):

(-1, 0) Yeah I just realized that :D

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

:)

OpenStudy (solomonzelman):

So, you just need the circles?

OpenStudy (anonymous):

Yes, I answered my question :D

OpenStudy (anonymous):

You helped me answer and understand it. Thank you very much, I appreciate it :)

OpenStudy (solomonzelman):

\(\large\color{black}{ \smile }\) you welcome!